### v1

# WaveNet’s Precision in EEG Classification

Abstract
This study introduces a WaveNet-based deep learning model designed to automate the classification of EEG signals into physiological, pathological, artifact, and noise categories. Traditional methods for EEG signal classification, which rely on expert visual review, are becoming increasingly impractical due to the growing complexity and volume of EEG recordings. Leveraging a publicly available annotated dataset from Mayo Clinic and St. Anne’s University Hospital, the WaveNet model was trained, validated, and tested on 209,232 samples with a 70/20/10 % split. The model achieved a classification accuracy exceeding previous CNN and LSTM-based approaches, and was benchmarked against a Temporal Convolutional Network (TCN) baseline. Notably, the model distinguishes noise and artifacts with high precision, although it reveals a modest but explainable degree of misclassification between physiological and pathological signals, reflecting inherent clinical overlap. WaveNet’s architecture, originally developed for raw audio synthesis, is well-suited for EEG data due to its use of dilated causal convolutions and residual connections, enabling it to capture both fine-grained and long-range temporal dependencies. The research also details the preprocessing pipeline, including dynamic dataset partitioning and normalization steps that support model generalization.
Index Terms:
EEG Signal Processing, Deep Learning, WaveNet Architecture, EEG Classification, Temporal Convolutional Networks, Neural Signal Analysis, artifact Detection.
I Introduction
The classification of EEG data into physiological, pathological, and noise categories is essential for accurate neurophysiological analysis and clinical diagnostics. Traditional methods rely heavily on expert visual review, which is not scalable given the increasing volume and complexity of EEG recordings. Trained EEG technologists are crucial in identifying abnormalities and enhancing patient outcomes. Manual EEG review by neurophysiologists and technologist is prone to errors, non-scalability, subjectivity, and inconsistencies due to varying expertise, fatigue, and cognitive biases [1]. In contrast, automated models [2, 3] are scalable, mitigate oversight, and use deep learning (DL) to discern subtle EEG patterns, improving reliability for precise diagnoses and treatments.
A Temporal Convolutional Network (TCN) is a type of DL neural network designed for processing sequential data. WaveNet is a specific type of TCN that excels at handling the intricate temporal characteristics of audio waveforms due to its architecture, which includes dilated causal convolutions and residual connections [4]. In WaveNet, causal convolutions ensure that the output at a time step depends on the current and previous time steps, mimicking the autoregressive nature of time series data like EEG signals. This is crucial as it prevents information leakage from future time steps into the current prediction.
EEG data are continuous signals reflecting the brain’s electrical activity, exhibiting complex temporal dependencies and dynamic variations [5]. The architecture is particularly well-suited for this task, as it can capture both fine-grained and long-range temporal dependencies in the data. This capability makes WaveNet a strong option for EEG data classification:
  • Temporal Dynamics: WaveNet’s ability to model long-range dependencies and temporal patterns aligns well with capturing EEG signals’ temporal dynamics. EEG signals fit this description perfectly, as they exhibit complex temporal dependencies and variations (see Figure 1 and 2 for examples).
  • Hierarchical Features: WaveNet’s hierarchical nature allows it to learn multi-scale features, crucial for identifying patterns in EEG data at different temporal resolutions. High-resolution features are captured by lower layers, while higher layers capture long-term, low-resolution patterns. This hierarchical feature extraction is crucial for modeling complex temporal dynamics in sequential data, which can allow the model to differentiate between physiological and pathological EEG data (Figures 1 and 2) and noise or artifacts.
  • High Resolution: WaveNet can capture detailed variations in EEG signals, essential for accurate classification, like the differences in the peaking behaviour of the EEG signals of pathological and healthy patients (Fig. 1 and Fig. 2).
Therefore, we propose using a WaveNet model to automate the classification of EEG data. To provide context, we also benchmark its performance against a TCN baseline. WaveNet’s architecture makes it well-suited for capturing the complex temporal patterns in EEG signals. Automating EEG data classification with WaveNet will not only enhance efficiency and accuracy but also ensure more consistent and reliable diagnostics, ultimately improving patient outcomes and advancing the field of neurophysiology.
Figure 1: Healthy physiological signals. Characterized by a steep incline and decline before and after the peak.
Figure 2: Example of abnormal brain activity observed in epilepsy-prone neurons is identified by a sharp wave transient. Indicated by the red arrow one can see this abnormal transient, a high frequency oscillation (HFO) can be observed riding atop the peak of the spikes.
II Related works
Accurate detection of epileptiform spikes is essential for the diagnosis and clinical management of epilepsy. Earlier methods relied heavily on manual annotation or rule-based algorithms, which are time consuming and subject to inter-rater variability, while recent deep learning approaches show promise in automating EEG spike classification by learning features directly from raw or minimally processed signals [6].
Initial work applied convolutional neural networks (CNNs) to this task. Johansen et al. [7] reported a CNN-based classifier with an area under the curve (AUC) of 95 %, effectively identifying epileptiform spikes with high sensitivity and low false positive rates. Nejedly et al. [3] extended this approach by combining graphoelement-derived features with both CNN and long short-term memory (LSTM) architectures, achieving F1 scores of 0.81 and 0.80, respectively. Their specialized patient tuned CNN model for pathological activity classification reached an average F1 score of 0.96. They also explored artifact rejection in intracerebral EEG using CNN-based classifiers [2], improving reliability by filtering out physiological noise from the input signals.
Subsequent research incorporated attention mechanisms into recurrent models to better capture temporal dependencies. Fukumori, Yoshida, and Tanaka [8] proposed a self-attentive recurrent neural network that dynamically emphasized relevant time steps, achieving an average accuracy of 90 % in epileptic spike detection. Similarly, Rukhsar and Tiwari [9] introduced an Attention Recurrent Neural Network (ARNN) that applied self-attention across multi-channel EEG data. Evaluated on datasets like CHB-MIT and canine EEG from the Mayo Clinic and North Carolina State University, the model achieved an average F1 score of 0.94 and outperformed LSTM and Vision Transformer baselines. However, ARNN remains a specialized model trained and evaluated within specific datasets, and its performance on cross-patient or cross-institution data remains an open question.
These developments reflect a broader trend toward hybrid deep learning architectures that combine spatial and temporal modeling. The integration of CNNs, RNNs, and attention mechanisms has improved classification accuracy and continues to shape efforts toward real-time and/or scalable EEG analysis.
III Dataset
The dataset for this article comprises annotated intracranial EEG recordings from Mayo Clinic (MN, USA) and St. Anne’s University Hospital (Brno, Czech Republic). The dataset includes EEG data labelled into three categories: physiological activity, pathological/epileptic activity, and artifactual signals. Each data sample of 3 seconds is stored in a .mat file with a sampling frequency of 5kHz and consists of a vector with dimensions $(1,15000)$ containing a total of 209,232 samples. Comprehensive instructions for dataset usage are provided in the associated scientific papers [2][3].
IV Methods
IV-A Preprocessing steps
The preprocessing steps involved loading EEG data from .mat files stored in directories, and organizing them into HDF5 format. This format optimizes storage and access speed while handling large datasets efficiently. After this preprocessing step, the data is dynamically split into distinct training, validation, and test sets. This splitting is achieved through random sampling of the data keys, ensuring separation and preventing data leakage between the sets during model evaluation and training. The entire pipeline is designed with mechanisms to maintain the integrity and separation of data throughout each stage of processing. Then, the input layer normalizes the input using Z-score normalization, adding $\varepsilon=1\times 10^{-8}$ to prevent division by zero.
IV-B Model structure
The WaveNet model is built using dilated causal convolutions to capture long-range dependencies in EEG signals. It employs seven sequential residual blocks with increasing dilation rates of $[1,2,4,8,16,32,64]$, allowing the receptive field to grow exponentially while maintaining the temporal ordering for time series modeling.
Each residual block includes:
  • Swish activation function defined as $\text{Swish}(x)=x\cdot\text{sigmoid}(x)$.
  • A dilated convolution layer with 32 filters and a kernel size of 3. For each convolutional layer with dilation rate $r$ and kernel $k$ size 3:
    $z_{i}(l)=\text{swish}\left(\sum_{j=0}^{k-1}w_{j}^{(l)}*x_{(i-j)}^{(l-1)}\right)$
  • A residual connection,
  • and A skip connection that feeds into a cumulative output stream.
These skip connections are summed and passed through two additional Conv1D layers. The first with 32 filters, and the second with 4 filters (equal to the number of output classes). All before being globally averaged across the time dimension via a GlobalAveragePooling1D layer. A final softmax activation layer produces the class probability distribution for each EEG segment.
To combat overfitting, an adaptive dropout mechanism is employed. This starts at an initial rate of 0.20 and dynamically adjusts throughout training based on a composite score combining validation accuracy, AUC, loss, and the relative gap between training and validation performance (for more detail see appendix A). The final observed dropout rate converged to 0.23 by the end of training.
Optimization is handled by the Adam [10] optimizer with an initial learning rate of 0.001, and L2 regularization with $\lambda=0.0001$ is applied to constrain model complexity and enhance generalization. To address class imbalance, particularly the dominance of physiological and pathological EEG patterns, the training pipeline applies dynamic class weighting called focal loss[11]. When using focal loss, class weights inversely proportional to class frequencies are passed as alpha parameters, focusing learning on underrepresented classes and down-weighting well-classified samples.
WaveNet models the conditional probability distribution over sequential data, estimating the joint probability of a sequence $P(x)=\prod_{t=1}^{T}P(x_{t}|x_{1},\dots,x_{t-1})$. This autoregressive structure makes might make WaveNet well-suited for modeling EEG signals, which exhibit strong temporal dependencies and nonstationary dynamics.
IV-C Baseline: Temporal Convolutional Network (TCN)
The Temporal Convolutional Network (TCN) was introduced as an additional baseline to provide a convolutional, non-autoregressive counterpart to WaveNet. While both architectures employ causal, dilated convolutions and residual connections, the TCN is tailored for classification rather than autoregressive synthesis. Its design kept receptive field and capacity comparable to WaveNet.
Architecture:
  • Residual stack of 8 causal Conv1D blocks with dilation rates $[1,2,4,8,16,32,64]$, kernel size $2$, and 8 filters per layer.
  • Each block contains two Conv1D layers with ReLU activations and weight normalization; residual paths use $1\times 1$ projections when needed.
  • Fixed dropout of $0.005$ within each block and Layer Normalization applied after residual addition.
  • Classification head: two Conv1D layers (8 then 4 filters), GlobalAveragePooling1D, and softmax output.
Training regimen:
  • Dataset: identical HDF5 source as WaveNet with 70/20/10 (train/validation/test) split.
  • Optimizer: Adam ($\text{lr}=1\times 10^{-3}$) with L2 regularization ($\lambda=1\times 10^{-4}$).
  • Loss: focal loss ($\gamma=2.0$) with class-weighted $\alpha$ parameters to correct imbalance.
  • Batch size: 16.
  • Early stopping on validation AUC (patience = 3) and model checkpointing based on macro-F1.
The baseline was implemented and trained using identical preprocessing and evaluation pipeline as WaveNet to ensure comparability. The core aim was to evaluate whether a purely convolutional, dilation-based classifier could rival WaveNet’s balance of temporal context modeling and residual learning.
IV-D Training summary
The wavenet model was trained on the Kaggle platform using a Tesla P100 GPU with optimized hardware acceleration. A custom focal loss [11] function addressed class imbalance, while adaptive dropout and adam were used to improve generalization. Key hyperparameters, including batch size, learning rate, and dilation settings, were carefully tuned for optimal performance. For a comprehensive overview of the training configuration and hardware specifics, see appendix A.
The TCN model training was conducted locally on a MSI Katana laptop using the NVIDIA GeForce RTX 3050 Laptop GPU with mixed-precision acceleration. A custom focal loss function [11] was used to mitigate class imbalance, while Adam optimization and a fixed dropout of 0.005 within residual blocks supported generalization. Key hyperparameters, including dilation rates, batch size (16), learning rate (1e-3), and number of filters (8), were tuned empirically for optimal performance.
V Results
The WaveNet model was trained on a curated dataset from St. Anne’s University Hospital and the Mayo clinic consisting of 292,232 EEG samples, split into 70% training, 20% validation, and 10% testing subsets. The model achieved an average F1 score of 0.94, outperforming prior methods such as the general CNN approach [2] and the graphoelements Convolutional LSTM model [3] which obtained an average F1 scores of 0.80 and 0.81 respectively (see Table III). While underperforming on the specialized model which reached a F1 of 0.96.
Training dynamics (Figure 3) demonstrate that the WaveNet reached  95% accuracy on both the training and validation set by the third epoch, stabilizing around 95%. The validation loss and accuracy curves (Figure 4) confirm stable convergence without significant overfitting.
The confusion matrix (Table I) illustrates WaveNet’s ability to discriminate noise and artifacts: noise was twice misclassified out of 3563 samples, and artifacts were seldom confused with pathological activity, but had some small recall issues with wrongly classifying physiological spikes. This is similar to the TCN baseline (Table II). It indicates that WaveNet and TCN reliably identify both transient and stationary non-neural signal components. However, some misclassification persisted between physiological and pathological EEG signals, which is consistent with the known clinical challenge of differentiating subtle abnormal brain activity from normal patterns.
TABLE I: Confusion Matrix for WaveNet Model Predictions on Test Set
True/Pred Noise Artifacts Physio. Patho.
Noise 3561 1 0 1
Artifacts 10 3372 371 47
Physiological 0 45 10866 175
Pathological 0 9 272 2166
TABLE II: Confusion Matrix for TCN Model Predictions on Test Set (default argmax threshold)
True/Pred Noise Artifacts Physio. Patho.
Noise 3542 43 5 0
Artifacts 1 3833 31 3
Physiological 0 657 9817 596
Pathological 0 47 205 2132
Figure 3: Accuracy figure, plotted every 150 batches each batch contains 32 samples for the training set. Yellow for the validation set plotted every epoch.
Figure 4: Validation and training loss figure, plotted every 150 batches each batch contains 64 samples for the training set. Yellow for the validation set plotted every epoch.
TABLE III: Classification Performance Metrics Across Different Models. LSTM Graphoelements model based on [3]. CNN General and CNN Specialized models based on [2].
Class/Metric CNN LSTM CNN WaveNet TCN
Graph. Gen. Spec.
Physiological
F1 0.86 0.90 0.98 0.96 0.93
Precision 0.85 0.93 0.97 0.94 0.98
Recall 0.87 0.87 0.97 0.98 0.89
Pathological
F1 0.73 0.64 0.90 0.90 0.83
Precision 0.66 0.56 0.90 0.96 0.78
Recall 0.82 0.74 0.91 0.89 0.89
Artifacts
F1 0.80 0.89 0.97 0.93 0.91
Precision 0.85 0.88 0.96 0.98 0.84
Recall 0.76 0.91 0.98 0.89 0.99
Noise
F1 0.98 $\sim$1.00 0.99
Precision 0.99 $\sim$1.00 $\sim$1.00
Recall 0.96 $\sim$1.00 $\sim$0.99
Macro avg.
F1 0.80 0.81 0.96 0.94 0.92
Precision 0.78 0.80 0.95 0.95 0.90
Recall 0.82 0.86 0.96 0.93 0.94
Overall, these results highlight the WaveNet model’s balance of sensitivity and specificity across classes, achieving precision and recall values exceeding 0.93 for most categories (Table III). This positions the model as a powerful tool for clinical EEG analysis, with clear advantages in suppressing irrelevant variance while preserving meaningful neurophysiological signals.
VI Conclusion
This study demonstrates that the WaveNet architectures, originally developed for audio synthesis, can be effectively repurposed for EEG classification. Its capacity to capture both fine-grained waveform subtleties and long-range temporal dependencies makes it well-suited for neurophysiological data. Leveraging a merged dataset from St. Anne’s University Hospital (Brno, Czech Republic) and the Mayo Clinic (Rochester, Minnesota, USA), the model achieved high validation and test accuracies, despite significant class imbalance and inter-subject variability.
Crucially, the use of dilated convolutions enabled the network to represent scale-variant features and rare event types, key characteristics of EEG and other biological signals. This temporal expressiveness, combined with adaptive regularization strategies such as dynamic dropout and focal loss with class reweighing, enhanced generalizability. The model’s strong cross-clinic performance suggests viability for real-world applications including seizure detection, artifact rejection, and closed-loop brain-computer interfaces.
WaveNet outperformed generalized CNN models across most metrics, while the specialized CNN achieved higher performance on physiological EEG signals (F1 0.96 vs. 0.98; Table III). The TCN baseline trained on the same preprocessing and dataset splits achieved performance comparable to WaveNet on several classes, demonstrating that a non-autoregressive convolutional architecture can capture EEG temporal patterns effectively. The specialized CNN reported by Nejedly et al. [2] highlights the trade off of highly tuned models: they maximize performance on particular signal types but require patient-specific tuning. WaveNet provides strong, consistent performance across classes, including improved pathological classification relative to the generalized CNN (F1 0.81 vs. 0.64). Which might reflect its deeper temporal receptive field and residual connections that capture long-range or irregular dynamics in EEG signals.
Notably, the WaveNet model was trained without patient specific tuning, yet achieved strong performance across clinically diverse EEG datasets. This outcome shows its capacity to generalize under real world conditions marked by variability and imbalance. Such inherent flexibility points toward a broader application in synthesize long range temporal dependencies and multi channel interactions in high density neural recordings. In particular, it opens the door to applying similar architectures to Neuropixels data within spike sorting pipelines like Kilosort, where waveform diversity and spatio-temporal structure require adaptable and scalable models. Building on these principles, future work will extend this approach to neuronal spike classification in large-scale electrophysiological recordings.
Spike-sorting approaches are shifting from fixed template clustering toward dynamic, interpretable, and biologically informed deep learning methods. Grounded in both neuroscience and machine learning theory, this framework aims to be scalable, transparent, and adaptable. Given WaveNet’s original design for raw audio synthesis, the model’s architecture could also be leveraged in future iterations to generate new healthy and epileptical spike, noise, and artifact signals for brain modeling input. This framework’s demonstrated generalizability positions it as a strong base model for hybrid systems, integrated with tools like Kilosort and extended via post hoc classifiers and explainability techniques. Future iterations will also incorporate tools such as Grad-CAM to elucidate model decision-making, enabling biologically meaningful interpretations of spike classifications.
Additionally, while the WaveNet model demonstrated cross-hospital generalization on diverse human EEG datasets, other architectures, such as the recently proposed Attention Recurrent Neural Network (ARNN) by Rukhsar and Tiwari [9], have reported impressive accuracy on long-duration, multi-channel seizure datasets, including CHB-MIT and canine EEG from the Mayo Clinic and North Carolina State University. While ARNN achieves strong accuracy, particularly in per-subject evaluations, it utilized 30 epochs for training, whereas our WaveNet model’s training had a hard stop at 10 epochs due to limited computing access. Future research should directly compare WaveNet and ARNN models on shared, heterogeneous EEG datasets, under similar training constraints, to evaluate their respective trade-offs in generalization, computational efficiency, and interpretability.
Both models leverage temporal pattern recognition for their high computational efficiency and accuracy. The ARNN model uses an LSTM-style recurrent gate to process sequences, while our WaveNet model employs dilated causal convolutions for temporal pattern recognition, followed by two convolutional ending layers. Given that traditional LSTM models, as baselines in other studies, have shown comparatively lower performance than more advanced temporal models on certain EEG datasets, an intriguing future direction is to explore hybrid architectures. Specifically, combining WaveNet’s unique capacity for capturing long-range temporal dependencies via dilated convolutions with the sequence modeling capabilities of recurrent networks could potentially enhance the efficiency and accuracy of models like ARNN.
In conclusion, this study demonstrates that WaveNet can serve as a powerful and generalizable framework for EEG analysis, capable of capturing fine grained waveform features while maintaining accuracy across heterogeneous datasets. Its capacity to synthesize long-range temporal dependencies and scale across diverse recording contexts highlights its utility not only for clinical applications such as seizure detection and artifact rejection, but also for broader neuroscientific research. Equally important, the architecture allows integration with interpretability methods, ensuring that future implementations can prioritize transparency and biological plausibility alongside performance. Together, these contributions establish WaveNet as a flexible and trustworthy foundation for advancing neural signal synthesis and classification, with potential extensions to large-scale electrophysiological data and hybrid deep learning approaches that further bridge neuroscience and machine learning.
Acknowledgment
I would like to thank the team at St. Anne’s University Hospital (Brno, Czech Republic) as well as the Mayo Clinic (Rochester, Minnesota, USA) for providing access to the EEG recordings that made this research possible. Special thanks go to Dr. Khubaib Ahmed for his valuable feedback on this article.
This research was conducted independently and without institutional funding. The author acknowledges the importance of interdisciplinary collaboration in bridging neuroscience and artificial intelligence, and hopes this work contributes to the development of more adaptive and interpretable neurotechnological tools.
Appendix A Detailed Training Configuration
  • Data Preparation: Batches were well-mixed, with our checks confirming that 0 % were highly clustered (defined as bigger than 90 % class dominance within a batch). This ensured effective global shuffling and diverse batches, as demonstrated by the representative analysis of the first five batches where the maximum class dominance ranged from 40.62 % to 68.75 %.
  • Loss Function: Custom Focal Loss with parameters gamma=2.0 and class-weight-derived alpha values to address class imbalance.
  • Hardware and Platform: Training was conducted on Kaggle using a Tesla P100-PCIE-16GB GPU (CUDA platform) with cuDNN version 90300 and XLA compilation for performance.
  • Model Architecture: WaveNet with 7 dilated causal convolutional layers, each with 32 filters, kernel size 3, Swish activation, and L2 regularization ($\lambda=0.0001$). Dilation rates: [1, 2, 4, 8, 16, 32, 64].
  • Training Duration: Total runtime was approximately 11 hours, 19 minutes, and 16 seconds.
  • Hyperparameters:
    • Input sequence length: 15000
    • Batch size: 32 (effective 64 via gradient accumulation)
    • Learning rate: 0.001
    • L2 regularization: 0.0001
  • Adaptive Dropout: The dropout rate was adjusted dynamically based on a composite score.
    Composite Score = (PrevAcc - ValAcc) $\times$ 1.5
    Composite Score = + (PrevAUC - AUC) $\times$ 0.8
    Composite Score = + (ValLoss - PrevLoss)
    Composite Score = + (RelGap - BaselineGap)
    Composite Score = + 25 // ValNoLearn
    Where:
    • PrevAcc: Validation accuracy from the previous epoch.
    • ValAcc: Current epoch’s validation accuracy.
    • PrevAUC: Validation AUC from the previous epoch.
    • AUC: Current epoch’s AUC.
    • ValLoss / PrevLoss: Validation loss from the current/previous epoch.
    • RelGap: $\frac{\max(\text{TrainAcc},10^{-8})-\text{ValAcc}}{\max(\text{TrainAcc},10^{-8})}$
    • BaselineGap: Constant (set to 0.02) defining an acceptable generalization gap.
    • ValNoLearn: $|\text{ValAcc}-\text{PrevAcc}|$ if less than 0.01, else 20.
    Positive Composite Scores indicate stagnation, performance degradation, or overfitting, which increase the dropout rate via(Fig. A1):
    New dropout = Current dropout + 0.05xComposite score
    This adaptive mechanism encourages regularization when learning stagnates or generalization degrades.
Figure A5: Dropout rate dynamics throughout training, governed by the composite performance score.